OpenStudy (anonymous):

Triangles ABC and LMN are given such that AB = 24, BC = 20, and AC = 40 LM = 6, MN = 5, and LN = (3a + 1) To be similar, which should be the value of a?

OpenStudy (anonymous):

6 5 4 3

OpenStudy (anonymous):

If AB/LM = BC/MN = AC/LN ... then they are similar AB/LM = 24/6 = 4 BC/MN = 20/5 = 4 AC / LN should also be 4 So , 40/(3a+1) = 4 40/4 = 3a+1 10 = 3a + 1 3a = 9 a = 3

OpenStudy (anonymous):

4th option

OpenStudy (anonymous):

Thanks! (again :P)